OpenStudy (unicwaan):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18, ...

OpenStudy (solomonzelman):

ok, your common ratio is>?

OpenStudy (unicwaan):

Okay

OpenStudy (solomonzelman):

your common ratio is not "okay" lol

OpenStudy (solomonzelman):

do you know what a common ratio is?

OpenStudy (unicwaan):

XD Oh I thought you were telling me that, I'm sorry. I do not know what a common ratio is.

OpenStudy (solomonzelman):

common ratio, is a number by which you multiply to get to the next term each time.

OpenStudy (unicwaan):

Oh okay so would it be (-1/6)?

OpenStudy (solomonzelman):

So, if I had 1, 2, 4, 8, 16 .... then my common ratio is 2 (and common ratio is denoted as r, so in this example r=2) --------------------------------------------- Yes, r=-1/6 is right

OpenStudy (unicwaan):

Okay! I undertsand that!

OpenStudy (solomonzelman):

Note: This is a geometric series (because it is multiplied times some number (referred to as common ratio (r) ).

OpenStudy (solomonzelman):

Do you know what it means for a series to converge and for a series to diverge?

OpenStudy (solomonzelman):

oh, for a sequence to diverge and converge

OpenStudy (unicwaan):

No I do not, that's what I'm coversed on. My lesson never went over it :(

OpenStudy (unicwaan):

confused**

OpenStudy (solomonzelman):

sequence, is just a list of terms. When you say that "sequence converges to 3" (for example 3) you are saying that as you take 1000th 100000th (and roughly speaking infinity-th terms) they all will be approximately 3 (and each time they will get closer and closer to 3)

OpenStudy (solomonzelman):

Say you started from 60, and you were to just divide by 1/6 (not -1/6). \(a_1=60\) \(a_2=10\) \(a_3=5/3\) \(a_4=5/18\) \(a_5=5/108\) and roughly speaking \(a_\infty=5/\infty=0\)

OpenStudy (unicwaan):

Ohhhhhh that makes sense okay

OpenStudy (solomonzelman):

No, you won't actually hit 0. Never! But you closer and closer approach to zero every time when you divide by 1/6. (this is the idea of a "limit" that you approach some value)

OpenStudy (unicwaan):

What about if something diverges?

OpenStudy (solomonzelman):

now, in our case we are multiplying (in the first post, should say multiplying too) by -1/6 so, all we change is that \(a_1=60\) \(a_2=-10\) \(a_3=5/3\) \(a_4=-5/18\) \(a_5=5/108\) and so on.... we still get \(a_\infty=\pm5/\infty=0\) (whether a positive or negative number is divided by infinity, you get 0 in either case)

OpenStudy (solomonzelman):

So the more terms you take, the close and close you approach what value?

OpenStudy (unicwaan):

0, since the number values get smaller and smaller

OpenStudy (solomonzelman):

yes, that is correct

OpenStudy (solomonzelman):

So that means that "sequence converges to zero"

OpenStudy (unicwaan):

^-^ Thanks a lot that makes sense!

OpenStudy (solomonzelman):

(and that is the only case - i.e. - sequence converges to zero - for a series to converge. If you want we can talk about the convergence of a series as well (later on in this thread))

OpenStudy (solomonzelman):

And you asked about divergence, i will answer that now

OpenStudy (unicwaan):

When I do reach series in my lesson I will come to you! and yes

OpenStudy (solomonzelman):

Say you got 1, 2, 4, 8, 16, 32 ....... what is the pattern here, can you tell me?

OpenStudy (unicwaan):

the next number is being multiplied by 2

OpenStudy (solomonzelman):

yes, r=2

OpenStudy (solomonzelman):

And as you take more and more terms you multiply times 2 more and more.... so you are going to go into infinity. (makes sense?)

OpenStudy (unicwaan):

Yes!

OpenStudy (solomonzelman):

that means that sequence diverges.

OpenStudy (unicwaan):

Oh okay so diverges means the increase (or decrease) infinitely while converges ia to get closer and closer to a certain number?

OpenStudy (solomonzelman):

so sequence can diverge in 2 cases. (1) if you have an alternating sequence that is not approaching zero. Such that \(A_n=4(-1)^n\), so your terms are going to be 4, -4, 4, -4, and so on.... and as n approaches infinity (n→∞) you don't know what your term is going to be because it can be either 4 or -4. (2) if your sequence goes into infinity (or negative infinity).

OpenStudy (solomonzelman):

yes, what you said is correct

OpenStudy (solomonzelman):

that is a good definition of convergence of a SEQUENCE.

OpenStudy (unicwaan):

Okay only for a sequence, got it.

OpenStudy (solomonzelman):

now, a series is the sum of terms in a sequence.

OpenStudy (solomonzelman):

think about it. Series is a sum of all terms in the sequence. So if sequence converges to 3 to -8 (or to anything besides 0) then you are going to be adding 3's -4's or something's forever and that will not have a defined sum (will go into ±∞)

OpenStudy (solomonzelman):

is this making some sense?

OpenStudy (unicwaan):

I believe so, yes

OpenStudy (solomonzelman):

yes, so for example in your case, your sequence converges to what?

OpenStudy (unicwaan):

The sequence converges to zero

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and the rule is that a geometric SERIES (a series that has all terms follow the pattern of multipling times some common ratio) will always converge if common ratio r is: -1>r>1

OpenStudy (unicwaan):

But in this case, the rule does not apply

OpenStudy (solomonzelman):

this rule always applies provided that ratio is: -1<r<1

OpenStudy (solomonzelman):

and so it is in your case, cuz r=-1/6 so it is between -1 and 1

OpenStudy (unicwaan):

Oh okay, I was thinking 0 was r, but r is the common ratio which was -1/6, That makes sense

OpenStudy (solomonzelman):

lets see, but why does the common ratio r has to satisfy -1<r<1? I will adress that now....

OpenStudy (solomonzelman):

for any geometric sequence you multiply times r to get to the next term. So lets say you start from \(a_1=5\) (or from any non-zero first term) \(\large\color{black}{ \displaystyle a_1=5}\) \(\large\color{black}{ \displaystyle a_2=5\cdot r}\) \(\large\color{black}{ \displaystyle a_3=5\cdot r^2}\) \(\large\color{black}{ \displaystyle a_4=5\cdot r^3}\) so on.... \(\large\color{black}{ \displaystyle a_n=5\cdot r^{n-1} }\)

OpenStudy (solomonzelman):

and then you get, roughly speaking \(\large\color{black}{ \displaystyle a_\infty=5\cdot r^{\infty -1}}\)

OpenStudy (unicwaan):

Makes sense

OpenStudy (solomonzelman):

so, if r\(\ge\)1 or if r\(\le\)-1 then the \(\large\color{black}{ \displaystyle r^{\infty -1}}\) part will go to either ∞ or -∞. but if -1<r<1 then you are going to have \(\large\color{black}{ \displaystyle r^{\infty -1}=0}\)

OpenStudy (solomonzelman):

your teacher will go over in detail, but the bottom line is that series is convergent, ONLY when sequence converges to 0. (there are many more tests for various series to converge/diverge that you will learn later in calc, but for geometric series, all you need to know is sequence converges to 0) and sequence convergence is when the \(a_\infty\) approaches one particular value.

OpenStudy (solomonzelman):

And for any geometric series \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ A_n}\) the sum is given by \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ A_n~=\frac{a_1}{1-{\rm r}}}\) in your case, the terms 60 is \(a_1\) and r=-1/6 so, \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ A_n~=\frac{60}{1-{\rm -\dfrac{1}{6}}}=\frac{60}{\dfrac{7}{6}}=360/7 }\)

OpenStudy (unicwaan):

:D Thank you for explaining so much to me! This really helped a lot. and I learn about that and summantion notation in my next lesson so thanks!

OpenStudy (solomonzelman):

Anytime!